OpenStudy (kaynick):

The population of Phoenix is about 1.728 x 10^6. The area of the city has an area of about 1.35 x 10^3 km^2. Find the number of people residing in Phoenix per square mile. (Population Density = Population / Area) Write your answer in standard form. Example: The standard form of 2.37 x 10^4 is 23,700

satellite73 (satellite73):

divide like it says \[\frac{1.728\times 10^6}{1.35\times 10^3}\]

satellite73 (satellite73):

if it was me, i would use a calculator